OpenStudy (chris215):

Find dy/dx if f(x) = (x + 1)^(2x)

OpenStudy (irishboy123):

take logs on both sides and see where you go :p use implicit differentiation

OpenStudy (anonymous):

\[let f(x)=y\] \[y=\left( x+1 \right)^{2x}\] ln y=2xln(x+1) diff

OpenStudy (chris215):

i got f '(x) = 2 (x+1)^(2 x-1) (x+(x+1) log(x+1))

OpenStudy (anonymous):

\[\frac{ 1 }{ y }\frac{ dy }{ dx }=2x \times \frac{ 1 }{ x+1 }+2 \ln \left( x+1 \right)\] \[\frac{ dy }{ dx }=y \left[ \frac{ 2x }{ x+1 }+2\ln \left( x+1 \right) \right]\] replace the value of y

zepdrix (zepdrix):

I got the same result Chris :) I'm pretty sure that's right.

OpenStudy (chris215):

thanks guys :)